OpenStudy (exoexoexo):

Janine has $12.50 in pennies(p), dimes(d), and half-dollars(h) in her pickle jar. She has 5 times as many dimes as half dollars and 5 times as many pennies as dimes a) if Janine has h half dollars, write algebraic expressions for the numbers of dimes and pennies that she has

jimthompson5910 (jim_thompson5910):

if `She has 5 times as many dimes as half dollars` then how many dimes does she have (in terms of h)?

OpenStudy (some.random.cool.kid):

..

OpenStudy (exoexoexo):

@jim_thompson5910 1/5d = h

jimthompson5910 (jim_thompson5910):

which is equivalent to d = 5h, good

jimthompson5910 (jim_thompson5910):

and if it says `5 times as many pennies as dimes` then what is an expression for the number of pennies (in terms of d dimes)?

OpenStudy (exoexoexo):

5p=d or 1/5p =d

jimthompson5910 (jim_thompson5910):

(# of pennies) = 5*(number of dimes) p = 5d

jimthompson5910 (jim_thompson5910):

so we know d = 5h p = 5d

jimthompson5910 (jim_thompson5910):

does that help you see how to finish?

OpenStudy (exoexoexo):

yes thank you

jimthompson5910 (jim_thompson5910):

you're welcome